OpenStudy (anonymous):

What are some multiples of 675? Above 15 x

OpenStudy (nowhereman):

The multiples of 675 are \[\{n\cdot 675 \mid n \in ℕ\}\]

OpenStudy (anonymous):

Wth

OpenStudy (nowhereman):

Sorry?

OpenStudy (anonymous):

I mean forexample, 15x 45 = 675

OpenStudy (nowhereman):

Do you mean factors?!?

OpenStudy (anonymous):

Yeah sorry

OpenStudy (nowhereman):

Well from that factorization you have there you can easily read the prime factors: \[675 = 3^5\cdot 5^2\] So all the possible factors of 675 are of the form \[3^i\cdot 5^j, i,j \in ℕ, i\leq 5, j \leq 2\]